OpenStudy (anonymous):

solve, using the addition and multiplication principles 5/11a - 1= 16/11

OpenStudy (anonymous):

ok \[5/11 * a - 1= 16/11\]

OpenStudy (anonymous):

so 1 = what in terms of a fraction with a denominator of 11?

OpenStudy (anonymous):

\[1 = 11/11\rightarrow 5/11*a - 11/11 = 16/11 \rightarrow \] add 11/11 to the left side to remove the -11/11 term, and add 11/11 to the right side because you have to to the same operation on each side to maintain equality here

OpenStudy (anonymous):

now \[5/11 * a - 11/11 + 11/11 = 16/11 + 11/11 \rightarrow 5/11*a = 27/11\]

OpenStudy (anonymous):

now using multiplication we can solve for a \[5/11*a = 27/11 \rightarrow (11/5)*(5/11)*a = (27/11)*(11/5)\] multiply both sides by 11/5 to reduce the coefficient on a to be equal to 1 so now we have \[a = (27/11) * (11/5) \rightarrow a = 27/5\]